# Edge densities of drawings of graphs with one forbidden cell

### v2

@@ 8 Drawing All Graphs without Some Cell Type
Proposition 8.1.
All connected simple graphs, except $K_{3}$, $K_{4}$ and the star graphs $K_{1,n}$ for $n\geq 0$, admit a simple drawing in which every cell is incident to at least one crossing.
Proof 8.2.
@@
Hence we may assume that $G\neq H$. Let $uv\in E(H)$ be an edge of $H$ and let $S$ be the set of vertices on the part of the circle $C$ bounded by (but not including) $u$ and $v$ that contains all other vertices of $H$. As $G\neq H$ and $G$ is connected, we may assume that $u$ has a neighbor $w$ with $w\notin S\cup\{u,v\}$. Let $R$ denote the part of the circle $C$ bounded by $u$ and $w$ that does not include $v$.
Let $\Gamma^{\prime}$ be the drawing that is obtained by moving all vertices in $S$ onto $R$ such that their clockwise order along $C$ is reversed relative to $\Gamma$. See Figure˜13 (top left) for an illustration. We argue that there are more crossings in $\Gamma^{\prime}$ than in $\Gamma$. Note that, since $H$ is an uncrossed cycle in $\Gamma$, all crossings in $\Gamma$ are either spanned by four vertices in $V\setminus S$ or by four vertices in $S\cup\{u,v\}$. All these crossings are also present in $\Gamma^{\prime}$, since the subdrawings of $\Gamma$ and $\Gamma^{\prime}$ on $V\setminus S$ and $S\cup\{u,v\}$, respectively, have the same vertex orders along $C$ (up to global reflection). Additionally, in $\Gamma^{\prime}$, the edge $vv^{\prime}\in E(H)$ with $v^{\prime}\neq u$ crosses $uw$. This contradicts crossing-maximality of $\Gamma$, so our assumption was faulty. Thus, all interior cells in $\Gamma$ are incident to a crossing.
Figure 13: Illustrations for Proposition˜8.1. Top left: Arguing that a crossing-maximal convex drawing has no interior uncrossed cells. Top right: Ensuring the outer cell is incident to a crossing in Case 1. Bottom: Arguing that the outer cell is incident to a crossing in Case 2.
@@
Case 1: Every edge of the convex hull of $V$ is an edge of $G$. Let $v_{0},\ldots,v_{n-1}$ be the vertices on $C$ in clockwise order. All subsequent index operations are implicitly performed modulo $n$. Let $v_{i}$ and $v_{j}$ be two non-adjacent vertices such that jij-i is minimal among all non-adjacent vertex pairs. Note that such a pair exists because $G$ is not complete and that the current case implies $j-i\geq 2$. Hence, placing $v_{j}$ in the space between $v_{i}$ and $v_{i+1}$ on $C$ gives a new drawing $\Gamma^{\prime}$; see Figure˜13 (top right) for a depiction. Since $v_{i}v_{j}\notin E$, the outer cell in $\Gamma^{\prime}$ is incident to a segment of the edge $v_{i}v_{i+1}$. Since further, $v_{i}v_{i+1}$ is crossed by $v_{j}v_{j+1}$ in $\Gamma^{\prime}$, this ensures that the outer cell is incident to a crossing. To finish Case 1, we are left with verifying that the move of $v_{j}$ between $\Gamma$ to $\Gamma^{\prime}$ did not introduce any uncrossed interior cells. For this, suppose there was an interior cell in $\Gamma^{\prime}$ bounded by a cycle $H\subseteq G$. The moved vertex $v_{j}$ cannot be in $H$, as all its incident edges except $v_{j}v_{i+1}$ are crossed by $v_{i}v_{i+1}$. Similarly, $v_{i+1},\ldots,v_{j-2}$ are not part of $H$ since $v_{i+1},\ldots,v_{j}$ form a convex drawing of a complete graph in $\Gamma^{\prime}$ (otherwise $v_{i}v_{j}$ was not a shortest missing edge). Further, all edges between $v_{j-1}$ and vertices $v_{j+2},\ldots,v_{i}$ except the edge $v_{j-1}v_{j+1}$ (if it exists) are crossed by the edge $v_{j}v_{j+1}$ in $\Gamma^{\prime}$, implying that, again, $v_{j-1}$ is not contained in $H$. Therefore, $H$ contains only vertices in $v_{j+1},v_{j+2},\ldots,v_{i}$. However, since $v_{j}$ does not change its position with respect to these vertices, $H$ also bounds a cell in $\Gamma$. This contradicts our assumption that all inner cells of $\Gamma$ have an incident crossing. Hence, $\Gamma^{\prime}$ is indeed a drawing of $G$ with the desired properties.
@@
Case 2: In $\Gamma$, two neighbors $u$ and $v$ on $C$ are not adjacent in $G$. Without loss of generality, let $u$ precede $v$ in clockwise order. Further, let $u^{\prime}$ (respectively $v^{\prime}$) be the neighbor of $u$ (respectively $v$) going clockwise (respectively counterclockwise) along the boundary of the outer cell of $\Gamma$. Since there are no crossings on the outer face of $\Gamma$, both $u^{\prime}$ and $v^{\prime}$ are cut-vertices in $G$, disconnecting $u$ and $v$ when removed. If $u^{\prime}\neq v^{\prime}$, then switching the positions of $u$ and $v$ introduces a new crossing between $uu^{\prime}$ and $vv^{\prime}$, while retaining all other crossings, see Figure˜13 (bottom left) — a contradiction to the crossing-maximality of $\Gamma$.
@@
So, we may assume that $u^{\prime}=v^{\prime}$. Moreover, by the previous reasoning, we may assume that any two non-adjacent vertices of $G$ that are neighbors on $C$ in $\Gamma$ are adjacent to a common cut-vertex. If both $u$ and $v$ have degree 1, then walking along $C$ eventually gives two non-adjacent neighboring vertices on $C$, both adjacent to $u^{\prime}$ and one of which of degree at least 2. This is guaranteed by the previous assumption and since $G$ is not a star. Without loss of generality, let $\deg(v)>1$. Then, switching the position of $u$ and $v$ on $C$ increases the number of crossings, since the edge $wv$ to any vertex $w\neq u^{\prime}$ adjacent to $v$ in $G$ now crosses $uu^{\prime}$. See Figure˜13 (bottom right). Once more, we reached a contradiction to crossing-maximality of $\Gamma$. Therefore, the outer cell of $\Gamma$ is also incident to a crossing and the claim is proven.
@@
With Proposition˜8.1 at hand, we can fully characterize the simple connected graphs that admit simple $\mathfrak{c}$-free drawing for every cell type $\mathfrak{c}$ whose boundary is not incident to any crossing.
@@
Let $\mathfrak{c}_{m}$ denote the cell type whose boundary is connected and bounded by $m$ vertices and $n$ uncrossed edge segments. Then, for every $m\geq 2$, every connected graph $G$ admits a simple $\mathfrak{c}_{m}$-free drawing except
  • $m=3$ and $G=K_{3}$,
  • for any $k\in\mathbb{N}$, $m=2k$ and $G$ the star on $2k+1$ vertices.
Proof 8.4.
Almost the entire statement is an immediate consequence of Proposition˜8.1. For the complete graph $K_{3}$ and star graphs, the unique (up to homeomorphism of the plane) simple drawing explains the exceptions. Finally, a planar drawing of $K_{4}$ contains only $\mathfrak{c}_{3}$-cells and a convex straight-line drawing of $K_{4}$ consists of a $\mathfrak{c}_{4}$-cell and four [Uncaptioned image]-cells, so any $\mathfrak{c}_{m}$ can be avoided by choosing one of the two drawings.
We remark that Proposition˜8.1 can be extended to disconnected graphs. In this case, if a graph $G$ contains two components with at least one edge each, there is a simple drawing of $G$ in which every cell is incident to a crossing. If otherwise, $G$ consists of a graph $H$ and some isolated vertices, $G$ admits such a drawing if and only if $H$ does. Corollary˜8.3 can be generalized to disconnected graphs in a similar way. For connected cell types without crossings, the only disconnected exception graph is $K_{3}+K_{1}$, which does not admit a $\mathfrak{c}_{3}$-free drawing. However, one might also want to consider cell types with disconnected boundary when including disconnected graphs. Then, the exception graphs are $K_{3}$, $K_{3}+K_{1}$ and $S+cK_{1}$ for some star graph $S$ and $c\geq 0$.
@@ 9 Concluding Remarks
The vast number of questions asked about other beyond planar graph classes – for example, about complexity of recognition and stretchability – can be carried over to our setting. However, our forbidden patterns rely on the topology of concrete drawings rather than only abstract combinatorics of the graph or the drawing (such as crossing edge pairs). Hence, the resulting graph classes are not always closed under taking subgraphs. In light of this, the results that almost all cell types can be avoided in drawings of $K_{n}$ from Section˜3 raise the question of which cell types can be avoided in drawings of any graph. For cell types whose boundary does not contain any crossing, we gave an answer in Section˜8 but a full characterization remains elusive.
Problem 9.3.

### v3

@@ 7 Drawing All Graphs without Some Cell Type
Theorem 7.1.
Let $G$ be a connected simple graph with $G\notin\{K_{3},K_{4}\}\cup\{K_{1,n}\mid n\geq 0\}$. Then $G$ admits a simple drawing in which every cell is incident to at least one crossing.
Proof 7.2.
@@
Hence we may assume that $G\neq H$. Thus, there is an edge $uv\in E(H)$ not on the convex hull of $V$. Let $S$ be the set of vertices on the part of the circle $C$ bounded by (but not including) $u$ and $v$ that contains all other vertices of $H$. As $G\neq H$ and $G$ is connected, we may assume that $u$ has a neighbor $w$ with $w\notin S\cup\{u,v\}$. Let $R$ denote the part of the circle $C$ bounded by $u$ and $w$ that does not include $v$.
Let $\Gamma^{\prime}$ be the drawing that is obtained by moving all vertices in $S$ onto $R$ such that their clockwise order along $C$ is reversed relative to $\Gamma$. See Figure˜12 (top left) for an illustration. We argue that there are more crossings in $\Gamma^{\prime}$ than in $\Gamma$. Note that, since $H$ is an uncrossed cycle in $\Gamma$, all crossings in $\Gamma$ are either spanned by four vertices in $V\setminus S$ or by four vertices in $S\cup\{u,v\}$. All these crossings are also present in $\Gamma^{\prime}$, since the subdrawings of $\Gamma$ and $\Gamma^{\prime}$ on $V\setminus S$ and $S\cup\{u,v\}$, respectively, have the same vertex orders along $C$ (up to global reflection). Additionally, in $\Gamma^{\prime}$, the edge $vv^{\prime}\in E(H)$ with $v^{\prime}\neq u$ crosses $uw$. This contradicts crossing-maximality of $\Gamma$, so our assumption was faulty. Thus, all interior cells in $\Gamma$ are incident to a crossing.
Figure 12: Illustrations for Theorem˜7.1. Top left: Arguing that a crossing-maximal convex drawing has no interior uncrossed cells. Top right: Ensuring the outer cell is incident to a crossing in Case 1. Bottom: Arguing that the outer cell is incident to a crossing in Case 2.
@@
Case 1: Every edge of the convex hull of $V$ is an edge of $G$. Let $v_{0},\ldots,v_{n-1}$ be the vertices on $C$ in cyclic order. All subsequent index operations are implicitly performed modulo $n$. Let $v_{i}$ and $v_{j}$ be two non-adjacent vertices such that jij-i is minimal among all non-adjacent vertex pairs. Note that such a pair exists because $G$ is not complete and that the current case implies $j-i\geq 2$. Hence, placing $v_{j}$ in the space between $v_{i}$ and $v_{i+1}$ on $C$ gives a new drawing $\Gamma^{\prime}$; see Figure˜12 (top right) for a depiction. Since $v_{i}v_{j}\notin E$, the outer cell in $\Gamma^{\prime}$ is incident to a segment of the edge $v_{i}v_{i+1}$. Since further $v_{i}v_{i+1}$ is crossed by $v_{j}v_{j+1}$ in $\Gamma^{\prime}$, this ensures that the outer cell is incident to a crossing. To finish Case 1, we are left with verifying that the move of $v_{j}$ between $\Gamma$ to $\Gamma^{\prime}$ did not introduce any uncrossed interior cells. For this, suppose there is an interior cell in $\Gamma^{\prime}$ bounded by a cycle $H\subseteq G$. The moved vertex $v_{j}$ cannot be in $H$, as all its incident edges except $v_{j}v_{i+1}$ are crossed by $v_{i}v_{i+1}$. Similarly, $v_{i+1},\ldots,v_{j-2}$ are not part of $H$ since $v_{i+1},\ldots,v_{j}$ form a convex drawing of a complete graph in $\Gamma^{\prime}$ (otherwise $v_{i}v_{j}$ was not a shortest missing edge). Further, all edges between $v_{j-1}$ and vertices $v_{j+2},\ldots,v_{i}$ except the edge $v_{j-1}v_{j+1}$ (if it exists) are crossed by the edge $v_{j}v_{j+1}$ in $\Gamma^{\prime}$, implying that, again, $v_{j-1}$ is not contained in $H$. Therefore, $H$ contains only vertices in $v_{j+1},v_{j+2},\ldots,v_{i}$. However, since $v_{j}$ does not change its position with respect to these vertices, $H$ also bounds a cell in $\Gamma$. This contradicts our assumption that all inner cells of $\Gamma$ have an incident crossing. Hence, $\Gamma^{\prime}$ is indeed a drawing of $G$ with the desired properties.
@@
Case 2: In $\Gamma$, two neighbors $u$ and $v$ on $C$ are not adjacent in $G$. Without loss of generality, let $u$ precede $v$ in clockwise order. Further, let $u^{\prime}$ and $v^{\prime}$ be the first neighbor along the boundary of the outer cell of $\Gamma$ starting clockwise from $u$ and counterclockwise from $v$, respectively . Since there are no crossings on the outer face of $\Gamma$, both $u^{\prime}$ and $v^{\prime}$ are cut-vertices in $G$, disconnecting $u$ and $v$ when removed. If $u^{\prime}\neq v^{\prime}$, then switching the positions of $u$ and $v$ introduces a new crossing between $uu^{\prime}$ and $vv^{\prime}$, while retaining all other crossings, see Figure˜12 (bottom left) — a contradiction to the crossing-maximality of $\Gamma$.
@@
So, we may assume that $u^{\prime}=v^{\prime}$. Moreover, by the previous reasoning, we may assume that any two non-adjacent vertices of $G$ that are neighbors on $C$ in $\Gamma$ are adjacent to a common cut-vertex. If both $u$ and $v$ have degree 1, then walking along $C$ eventually gives two non-adjacent neighboring vertices on $C$, both adjacent to $u^{\prime}$ and one of which of degree at least 2. This is guaranteed by the previous assumption and since $G$ is not a star. Without loss of generality, let $\deg(v)>1$. Then, switching the position of $u$ and $v$ on $C$ increases the number of crossings, since the edge $wv$ to any vertex $w\neq u^{\prime}$ adjacent to $v$ in $G$ now crosses $uu^{\prime}$. See Figure˜12 (bottom right). Once more, we reached a contradiction to crossing-maximality of $\Gamma$. Therefore, the outer cell of $\Gamma$ is also incident to a crossing and the claim is proven.
@@
With Theorem˜7.1 at hand, we can fully characterize the simple connected graphs that admit a simple $\mathfrak{c}$-free drawing for every (non-degenerate) cell type $\mathfrak{c}$ whose boundary is not incident to any crossing.
@@
For an integer $m\geq 2$, let $\mathfrak{c}_{m}$ denote the cell type whose boundary is connected and bounded by $m$ vertices and $m$ uncrossed edge segments. Let $G$ be a connected simple graph. Then $G$ admits a simple $\mathfrak{c}_{m}$-free drawing, unless
  • $m=3$ and $G=K_{3}$, or
  • $m$ is even and $G=K_{1,m/2}$.
Proof 7.4.
For all connected simple graphs $G$ with $G\notin\{K_{3},K_{4}\}\cup\{K_{1,k}\mid k\geq 0\}$, the statement is an immediate consequence of Theorem˜7.1. For $K_{4}$, a plane drawing contains only $\mathfrak{c}_{3}$-cells and a convex straight-line drawing consists of a $\mathfrak{c}_{4}$-cell and four [Uncaptioned image]-cells, so any $\mathfrak{c}_{m}$ can be avoided by choosing one of the two drawings. The unique drawing of $K_{1,0}=K_{1}$ has no cell of type $\mathfrak{c}_{m}$ for any $m\geq 2$. Finally, for the complete graph $K_{3}$ and the star graphs $K_{1,k}$, the unique (up to homeomorphism of the plane) simple drawings contain only the cell type $\mathfrak{c}_{3}$ and $\mathfrak{c}_{2k}$, respectively, which both explains the exceptions and shows the statement also for these graphs.
We remark that Theorem˜7.1 can be extended to disconnected graphs. In this case, if a graph $G$ contains two components with at least one edge each, there is a simple drawing of $G$ in which every cell is incident to a crossing. If otherwise, $G$ consists of a graph $H$ and some isolated vertices, $G$ admits such a drawing if and only if $H$ does.
Corollary˜7.3 can be generalized to disconnected graphs in a similar way. For connected cell types without crossings, the only disconnected exception graph is $K_{3}+K_{1}$, which does not admit a $\mathfrak{c}_{3}$-free drawing. However, one might also want to consider cell types with disconnected boundary and no incident crossing when including disconnected graphs. Then, the only exception graphs are $K_{3}$, $K_{3}+K_{1}$ and $S+nK_{1}$ for any star graph $S$ and $n\geq 0$, with respect to the cells that appear in their unique drawings.
@@ 8 Concluding Remarks
The vast number of questions asked about other beyond planar graph classes – for example, about complexity of recognition and stretchability – can be carried over to our setting. However, our forbidden patterns rely on the topology of concrete drawings rather than only abstract combinatorics of the graph or the drawing (such as crossing edge pairs). Hence, the resulting graph classes are not always closed under taking subgraphs. In light of this, the results that almost all cell types can be avoided in drawings of $K_{n}$ from Section˜3 raise the question of which cell types can be avoided in drawings of any graph. For cell types whose boundary does not contain any crossing, we gave an answer in Section˜7 but a full characterization remains elusive.
Problem 8.3.
@@
Apart from our contributions about forbidden cells, we also improved the lower bound on the edge density of simple graphs that admit a quasiplanar drawings from $7n-\mathcal{O}(1)$ to $7.5n-\mathcal{O}(1)$. However, a linear gap remains to the upper bound of 8n208n-20.
Problem 8.4.
Are there quasiplanar drawings of $n$-vertex simple graphs with $8n-\mathcal{O}(1)$ edges?
Acknowledgements
We thank the various anonymous referees, who, through their valuable comments, helped to significantly improve the presentation of this paper.